what is the value of K for 2x+ky=5 if this
equation is perpendicular to 3x+y=-2?

Answer by uma(370)

2x + ky = 5
This is the equation of a line.
Slope of this line = -2/k
The other line is given to be 3x + y = - 2
==> slope of this line = -3
As the lines are perpendicular, the product of their slopes = - 1
==> (-2/k)(-3) = - 1
==> 6/k = -1
==> 6 = - k
==> 6/-1 = -k/-1
==> - 6 = k
Thus k = - 6

Good Luck!!!